# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ D.1 Setup
We draw our prompts (prefixes) from a selection of books in Books3; we verify extraction against the corresponding suffixes drawn from Books3. Books3 is a common pretraining dataset that is known to have been used to train several open-weight LLMs (Appendix C.1, [104, 76]. (See models, below). We list the $50$ selected books in Table LABEL:app:tab:sliding-books.We began our project with The Great Gatsby [46]. This choice stems from a conversation in 2023 while writing a different paper [76], when Anthropic announced its 100,000100{,}000-token context for Claude with the example of prompting with the entirety of The Great Gatsby [4]. We then continued our selection process with the books listed with the associated plaintiffs in the (amended) class action complaint of Kadrey et al. v. Meta, Inc. [70, pp. 4-5]. This is how we sourced the first $13$ in-copyright books from Books3. We expanded to additional books by these plaintiffs, and also added in some generally popular books that we chose among the project team (e.g., Rowling [106], Tolkien [126], Camus [13], Heller [61]) and some (less publicly popular) academic books based on our personal preferences (e.g., Zittrain [139], Barolini [7]). (This team, after all, is composed of academics.) We deliberately include selections from the public domain (Carroll [19], Fitzgerald [46], Joyce [68]), and an example of a book that was published under a permissive CC license (Doctorow [36]).We added in Carroll [19] because of concurrent work that was published on arXiv just as we posted our own [83]. To round out the list, we also selected books at random from the manifest file that accompanies the Books3 archive (e.g., Jacobs [67], Rouighi [105], Cogburn [25]). Once we observed that certain very popular books were highly memorized by Llama 3.1 70B, we added additional bestsellers to our set (e.g., Martin [85], Brown [11], Meyer [87], Orwell [100], Sandberg [112]).
@@
Altogether, we run this set of experiments on $17$ completion-style (i.e., non-chatbot) models. We includes results for $10$ of these models: Pythia 12B [9], Llama 1 13B, Llama 1 65B [127], Llama 2 70B [128], Llama 3 70B, Llama 3.1 70B [54], DeepSeek v1 67B [35], Qwen 2.5 72B [103], Gemma 2 27B [121], and Phi 4 (a 14B model) [2]. For brevity, the $4$ (smaller) models we do not include in our results are Pythia 6.7B, Llama 1 7B, Llama 2 7B, Llama 2 13B, Llama 3 8B, Llama 3.1 8B, and DeepSeek v1 7B. We note that Llama 3 8B and Llama 3.1 8B exhibit an enormous degree of memorization on some books, for models in this size class (e.g., 1984).We load all models with float16—except for Llama 3, Llama 3.1, and Qwen 2.5 models. These models were explicitly trained to work with bfloat16. In an earlier version of this work, we reported results for Llama 3.1 that used float16. Slight differences in results in this version are attributed to the change to bfloat16. (We have otherwise confirmed our results are deterministic, with respect to the hardware that we use throughout.) Regardless, the overall results are similar, and the conclusions remain the same. We picked these models because they each fit (with respect to $16$-bit format) on $4$ A100s, and because they fall into three different categories that we wanted to investigate: (1) models that we know with certainty were trained on Books3, (2) models where we have reason to believe were trained on copyrighted books (whether drawn directly from Books3 or from other sources), and (3) a model that we trust was not trained on whole copyrighted books. For the first category, when we generate verbatim completions for these models on Books3 data, we can very reasonably conclude that we have extracted data that these models have memorized during training. We know that these data were members of the training data; extraction success (and thus memorization) is a standard conclusion in this setting [e.g., 9, 60, 101]. For the second category, we would like to be able to do valid membership inference of copyrighted data. The third category is also to help confirm validity; as a baseline, we intentionally include a model for which we do not expect to be able to extract much (if any) copyrighted books data.
@@
    Category 1: Pythia 12B, Llama 1 13B, Llama 1 65B, Llama 2 70B, Llama 3 70B, and Llama 3.1 70B. We know that Pythia and Llama models were training on Books3. Pythia was trained on the Pile [47] (which contains Books3), based on EleutherAI’s extensive documentation. We include only the largest ($12$-billion parameter) Pythia model.Llama 1 and Llama 2 are known with absolute certainty to have been trained on Books3. The Llama 1 report explicitly says so. It came out in discovery during Kadrey et al. v. Meta, Inc. [71] that Llama 2 models were all also definitively trained on Books3. Both in documents in discovery and Meta’s otherwise public documentation, it is repeatedly noted that the development of Llama 3 continued to use the same data practices as prior versions. For minor-versioned Llama 3 models (like Llama 3.1), it is similarly documented by Meta in release reports that the same internal workflows were used, including for dataset curation [54]. While this is not as definitive evidence that these models were trained on Books3, based on what is public information about the training process and prior Llama models, it is widely believed in the ML community that Llama 3 (and associated minor versions) models were trained on Books3 (as well as additional books data from LibGen). We consider this sufficient to conclude with confidence that Llama 3 and Llama 3.1 were trained on Books3: Books3 text sequences were members of the training dataset. (And regardless of this reasoning, our validity experiments enable us to independently conclude this. See Section 5 and Appendix E.)
@@
    Category 3: Phi 4. For validity reasons, we also want to compare on a model that definitively should not have been trained on Books3. (See also Appendix E.) We pick Phi 4 for these reasons, as it is a very high quality for its size class and Abdin et al. [2] note in the Phi 4 report that it was trained predominantly on synthetic data:
    • Synthetic data constitutes the bulk of the training data for phi-4 and is generated using a diverse array of techniques, including multi-agent prompting, self-revision workflows, and instruction reversal. …[Phi 4 was also trained on] web content, licensed books, and code repositories to extract seeds for the synthetic data pipeline. Abdin et al. [2] (emphasis added)
    Phi 4 is similar in size to Pythia 12B and Llama 1 13B, which facilitates cross-model comparisons of extraction risk. (This is because the amount of memorization, reasonably, also varies according to model size; see also Carlini et al. [18].)
@@
Table 3: The $50$ books drawn from Books3 that we evaluated with our sliding-window probabilistic extraction procedure. See Sections 4 and 6, as well as Appendices D.1D.2.
@@
1 Margaret Atwood The Handmaid’s Tale [5] 1985 © the-eye.eu/public/Books/Bibliotik/M/Margaret Atwood (1985) The Handmaid_s Tale [retail].epub.txt
@@
2 Teodolinda Barolini Dante and the Origins of Italian Literary Culture [7] 2006 © the-eye.eu/public/Books/Bibliotik/D/Dante and the Origins of Italia - Barolini, Teodolinda;.epub.txt
@@
3 Dan Brown The Da Vinci Code [11] 2003 © the-eye.eu/public/Books/Bibliotik/D/Dan Brown - The Da Vinci Code.epub.txt
@@
4 Albert Camus (Justin O’Brien, translator) The Myth of Sisyphus [13] 1955 © the-eye.eu/public/Books/Bibliotik/2/2013(orig1942) Albert Camus - The Myth of Sisyphus[Transl Justin O’Brien]_Ral.epub.txt
@@
5 Lewis Carroll Alice’s Adventures in Wonderland [19] 1865 PD the-eye.eu/public/Books/Bibliotik/A/Alice’s Adventures in Wonderlan - Lewis Carroll.epub.txt
@@
6 Ta-Nehisi Coates The Beautiful Struggle [22] 2009 © the-eye.eu/public/Books/Bibliotik/T/The Beautiful Struggle - Ta-Nehisi Coates.epub.txt
@@
7 Ta-Nehisi Coates We Were Eight Years in Power: An American Tragedy [23] 2017 © the-eye.eu/public/Books/Bibliotik/W/We Were Eight Years in Power - Ta-Nehisi Coates.epub.txt
@@
8 Ta-Nehisi Coates The Water Dancer: A Novel [24] 2019 © the-eye.eu/public/Books/Bibliotik/T/The Water Dancer - Ta-Nehisi Coates;.epub.txt
@@
9 Jon Cogburn Dungeons and Dragons and Philosophy [25] 2012 © the-eye.eu/public/Books/Bibliotik/2/2012 Jon Cogburn - Dungeons and Dragons and Philosophy - Raiding the Temple of Wisdom_Rsnl.epub.txt
10 Junot Díaz Drown [38] 1996 © the-eye.eu/public/Books/Bibliotik/J/Junot Diaz - Drown.epub.txt
@@
11 Junot Díaz The Brief Wondrous Life of Oscar Wao [39] 2007 © the-eye.eu/public/Books/Bibliotik/T/The Brief Wondrous Life of Oscar Wao.epub.txt
@@
12 Junot Díaz This Is How You Lose Her [40] 2012 © the-eye.eu/public/Books/Bibliotik/T/This Is How You Lose Her-Diaz.epub.txt
@@
13 Cory Doctorow Down and Out in the Magic Kingdom [36] 2004 $C$-BY-NC-SA the-eye.eu/public/Books/Bibliotik/D/Down and Out in the Magic Kingd - Cory Doctorow.epub.txt
@@
14 Carol Ann Duffy The World’s Wife [37] 2001 © the-eye.eu/public/Books/Bibliotik/T/The World_s Wife - Carol Ann Duffy.epub.txt
@@
15 Jennifer Egan A Visit from the Goon Squad [41] 2010 © the-eye.eu/public/Books/Bibliotik/A/A\ Visit\ from\ the\ Goon\ Squad\ -\ Jennifer\ Egan.epub.txt
@@
16 Christopher Farnsworth The President’s Vampire [43] 2011 © the-eye.eu/public/Books/Bibliotik/T/The President’s Vampire - Christopher Farnsworth.epub.txt
@@
17 F. Scott Fitzgerald The Great Gatsby [46] 1925 PD the-eye.eu/public/Books/Bibliotik/T/The Great Gatsby - F. Scott Fitzgerald.epub.txt
@@
18 Malcom Gladwell Blink: The Power of Thinking Without Thinking [49] 2005 © the-eye.eu/public/Books/Bibliotik/B/Blink - Malcolm Gladwell.epub.txt
@@
19 Christopher Golden Dead Ringer [50] 2016 © the-eye.eu/public/Books/Bibliotik/C/Christopher Golden - Dead Ringers [retail].epub.txt
20 Christopher Golden Ararat [51] 2017 © the-eye.eu/public/Books/Bibliotik/A/Ararat - Christopher Golden.epub.txt
@@
21 Andrew Sean Greer The Confessions of Max Tivoli [55] 2005 © the-eye.eu/public/Books/Bibliotik/T/The Confessions of Max Tivoli - Andrew Sean Greer.epub.txt
@@
22 John Grisham Theodore Boone: The Fugitive [56] 2015 © the-eye.eu/public/Books/Bibliotik/T/The Fugitive - John Grisham.epub.txt
@@
23 Mark Haddon The Curious Incident of the Dog in the Night-Time [57] 2003 © the-eye.eu/public/Books/Bibliotik/T/The Curious Incident of the Dog in the Night-Time - Mark Haddon.epub(1).txt
24 Joseph Heller Catch-22 [61] 2018 © the-eye.eu/public/Books/Bibliotik/C/Catch-22 - Joseph Heller.epub.txt
@@
25 David Henry Hwang M. Butterfly [65] 1988 © the-eye.eu/public/Books/Bibliotik/M/M. Butterfly - David Henry Hwang.epub.txt
26 Betty E.M. Jacobs All the Onions [67] 1999 © the-eye.eu/public/Books/Bibliotik/A/All the Onions - Betty E. M. Jacobs.epub.txt
@@
27 James Joyce Ulysses [68] 1922 PD the-eye.eu/public/Books/Bibliotik/U/Ulysses - James Joyce - Penguin Group - 2000.epub.txt
@@
28 Richard Kadrey Sandman Slim [69] 2009 © the-eye.eu/public/Books/Bibliotik/S/Sandman Slim - Richard Kadrey.epub.txt
29 Matthew Klam Who Is Rich? [72] 2017 © the-eye.eu/public/Books/Bibliotik/W/Who Is Rich_ - Matthew Klam.epub.txt
@@
30 Laura Lippman After I’m Gone [80] 2014 © the-eye.eu/public/Books/Bibliotik/A/After I’m Gone - Laura Lippman.epub.txt
31 Laura Lippman Sunburn [81] 2018 © the-eye.eu/public/Books/Bibliotik/S/Sunburn - Laura Lippman.epub.txt
@@
32 George R.R. Martin A Game of Thrones [85] 1996 © the-eye.eu/public/Books/Bibliotik/A/A Game of Thrones - George R. R. Martin.epub.txt
33 Stephenie Meyer Twilight [87] 2005 © the-eye.eu/public/Books/Bibliotik/T/Twilight - Stephenie Meyer.epub.txt
34 Toni Morrison Beloved [89] 1987 © the-eye.eu/public/Books/Bibliotik/B/Beloved - Toni Morrison.epub.txt
@@
35 Yōko Ogawa The Memory Police [96] 2019 © the-eye.eu/public/Books/Bibliotik/T/The Memory Police - Yoko Ogawa.epub.txt
@@
37 Philip Pullman The Subtle Knife [102] 1997 © the-eye.eu/public/Books/Bibliotik/P/Pullman - The Subtle Knife.epub.txt
38 Ramzi Rouighi The Making of a Mediterranean Emirate [105] 2011 © the-eye.eu/public/Books/Bibliotik/T/The Making of a Mediterranean E - Rouighi, Ramzi;.epub.txt
39 J.K. Rowling Harry Potter and the Sorcerer’s Stone [106] 1998 © the-eye.eu/public/Books/Bibliotik/H/Harry_Potter_ and_the_Sorcerers_Stone-Rowling.epub.txt
40 J.K. Rowling Harry Potter and the Goblet of Fire [107] 2000 © the-eye.eu/public/Books/Bibliotik/H/Harry_Potter_ and_the_Goblet_of_Fire-Rowling.epub.txt
41 J.D. Salinger The Catcher in the Rye [110] 1951 © the-eye.eu/public/Books/Bibliotik/T/The Catcher in the Rye - J. D. Salinger.epub.txt
42 Sheryl Sandberg Lean In [112] 2013 © the-eye.eu/public/Books/Bibliotik/L/Lean In - Sheryl Sandberg.epub.txt
43 Sarah Silverman The Bedwetter [117] 2010 © the-eye.eu/public/Books/Bibliotik/T/The Bedwetter - Sarah Silverman.epub.txt
44 Rachel Louise Snyder No Visible Bruises [118] 2019 © the-eye.eu/public/Books/Bibliotik/N/No Visible Bruises - Rachel Louise Snyder.epub.txt
45 Lysa TerKeurst Unglued [122] 2012 © the-eye.eu/public/Books/Bibliotik/L/Lysa TerKeurst, Unglued.epub.txt
46 Lysa TerKeurst Embraced [123] 2018 © the-eye.eu/public/Books/Bibliotik/E/Embraced - Lysa Terkeurst.epub.txt
47 J.R.R. Tolkien The Hobbit [126] 1937 © the-eye.eu/public/Books/Bibliotik/T/The Hobbit (Houghton Mifflin Harcourt) (75th Anniversary Edition) [Epub] - J.R.R. Tolkien.epub.txt
48 Jacqueline Woodson Brown Girl Dreaming [134] 2014 © the-eye.eu/public/Books/Bibliotik/B/Brown Girl Dreaming - Jacqueline Woodson.epub.txt
49 Jacqueline Woodson Another Brooklyn [135] 2016 © the-eye.eu/public/Books/Bibliotik/A/Another Brooklyn - Jacqueline Woodson.epub.txt
50 Jonathan Zittrain The Future of the Internet and How to Stop It [139] 2008 © the-eye.eu/public/Books/Bibliotik/J/Jonathan Zittrain - The Future of the Internet.epub.txt

### v4

@@ D.1 Setup
We draw our prompts (prefixes) from a selection of books in Books3; we verify extraction against the corresponding suffixes drawn from Books3. Books3 is a common pretraining dataset that is known to have been used to train several open-weight LLMs (Appendix C.1, [reisnerbooks3, lee2023talkin]. (See models, below). We list the $50$ selected books in Table LABEL:app:tab:sliding-books.
We began our project with The Great Gatsby [gatsby]. This choice stems from a conversation in 2023 while writing a different paper [lee2023talkin], when Anthropic announced its 100,000100{,}000-token context for Claude with the example of prompting with the entirety of The Great Gatsby [claudegatsby]. We then continued our selection process with the books listed with the associated plaintiffs in the (amended) class action complaint of Kadrey et al. v. Meta, Inc. [kadreyamendedconsolidated, pp. 4-5]. This is how we sourced the first $13$ in-copyright books from Books3. We expanded to additional books by these plaintiffs, and also added in some generally popular books that we chose among the project team (e.g., hp1, hobbit, sisyphus, catch22) and some (less publicly popular) academic books based on our personal preferences (e.g., zittrain, dante). (This team, after all, is composed of academics.) We deliberately include selections from the public domain (alice, gatsby, ulysses), and an example of a book that was published under a permissive CC license (downandout).We added in alice because of concurrent work that was published on arXiv just as we posted our own [ma2025memorizationcloselookbooks]. To round out the list, we also selected books at random from the manifest file that accompanies the Books3 archive (e.g., onions, emirate, dungeons). Once we observed that certain very popular books were highly memorized by Llama 3.1 70B, we added additional bestsellers to our set (e.g., thrones, davinci, twilight, 1984, leanin).
@@
Altogether, we run this set of experiments on $17$ completion-style (i.e., non-chatbot) models. We includes results for $10$ of these models: Pythia 12B [biderman2023pythia], Llama 1 13B, Llama 1 65B [touvron2023llamaopenefficientfoundation], Llama 2 70B [llama2], Llama 3 70B, Llama 3.1 70B [llama3], DeepSeek v1 67B [deepseekv1], Qwen 2.5 72B [qwen2.5], Gemma 2 27B [gemma2], and Phi 4 (a 14B model) [phi4]. For brevity, the $4$ (smaller) models we do not include in our results are Pythia 6.7B, Llama 1 7B, Llama 2 7B, Llama 2 13B, Llama 3 8B, Llama 3.1 8B, and DeepSeek v1 7B. We note that Llama 3 8B and Llama 3.1 8B exhibit an enormous degree of memorization on some books, for models in this size class (e.g., 1984).
We load all models with float16—except for Llama 3, Llama 3.1, and Qwen 2.5 models. These models were explicitly trained to work with bfloat16. In an earlier version of this work, we reported results for Llama 3.1 that used float16. Slight differences in results in this version are attributed to the change to bfloat16. (We have otherwise confirmed our results are deterministic, with respect to the hardware that we use throughout.) Regardless, the overall results are similar, and the conclusions remain the same.
We picked these models because they each fit (with respect to $16$-bit format) on $4$ A100s, and because they fall into three different categories that we wanted to investigate: (1) models that we know with certainty were trained on Books3, (2) models where we have reason to believe were trained on copyrighted books (whether drawn directly from Books3 or from other sources), and (3) a model that we trust was not trained on whole copyrighted books.
For the first category, when we generate verbatim completions for these models on Books3 data, we can very reasonably conclude that we have extracted data that these models have memorized during training. We know that these data were members of the training data; extraction success (and thus memorization) is a standard conclusion in this setting [e.g., biderman2023pythia, hayes2025measuringmemorizationlanguagemodels, prashanth2024recite]. For the second category, we would like to be able to do valid membership inference of copyrighted data. The third category is also to help confirm validity; as a baseline, we intentionally include a model for which we do not expect to be able to extract much (if any) copyrighted books data.
@@
    Category 1: Pythia 12B, Llama 1 13B, Llama 1 65B, Llama 2 70B, Llama 3 70B, and Llama 3.1 70B. We know that Pythia and Llama models were training on Books3. Pythia was trained on the Pile [gao2020pile] (which contains Books3), based on EleutherAI’s extensive documentation. We include only the largest ($12$-billion parameter) Pythia model.
    Llama 1 and Llama 2 are known with absolute certainty to have been trained on Books3. The Llama 1 report explicitly says so. It came out in discovery during Kadrey et al. v. Meta, Inc. [kadrey] that Llama 2 models were all also definitively trained on Books3.
    Both in documents in discovery and Meta’s otherwise public documentation, it is repeatedly noted that the development of Llama 3 continued to use the same data practices as prior versions. For minor-versioned Llama 3 models (like Llama 3.1), it is similarly documented by Meta in release reports that the same internal workflows were used, including for dataset curation [llama3]. While this is not as definitive evidence that these models were trained on Books3, based on what is public information about the training process and prior Llama models, it is widely believed in the ML community that Llama 3 (and associated minor versions) models were trained on Books3 (as well as additional books data from LibGen). We consider this sufficient to conclude with confidence that Llama 3 and Llama 3.1 were trained on Books3: Books3 text sequences were members of the training dataset. (And regardless of this reasoning, our validity experiments enable us to independently conclude this. See Section 5 and Appendix E.)
@@
    Category 3: Phi 4. For validity reasons, we also want to compare on a model that definitively should not have been trained on Books3. (See also Appendix E.) We pick Phi 4 for these reasons, as it is a very high quality for its size class and phi4 note in the Phi 4 report that it was trained predominantly on synthetic data:
    • Synthetic data constitutes the bulk of the training data for phi-4 and is generated using a diverse array of techniques, including multi-agent prompting, self-revision workflows, and instruction reversal. …[Phi 4 was also trained on] web content, licensed books, and code repositories to extract seeds for the synthetic data pipeline. phi4 (emphasis added)
    Phi 4 is similar in size to Pythia 12B and Llama 1 13B, which facilitates cross-model comparisons of extraction risk. (This is because the amount of memorization, reasonably, also varies according to model size; see also carlini2023quantifying.)
@@
Table 3: The $50$ books drawn from Books3 that we evaluated with our sliding-window probabilistic extraction procedure. See Sections 4 and 6, as well as Appendices D.1 & D.2.
@@
1 Margaret Atwood The Handmaid’s Tale [handmaid] 1985 © the-eye.eu/public/Books/Bibliotik/M/Margaret Atwood (1985) The Handmaid_s Tale [retail].epub.txt
@@
2 Teodolinda Barolini Dante and the Origins of Italian Literary Culture [dante] 2006 © the-eye.eu/public/Books/Bibliotik/D/Dante and the Origins of Italia - Barolini, Teodolinda;.epub.txt
@@
3 Dan Brown The Da Vinci Code [davinci] 2003 © the-eye.eu/public/Books/Bibliotik/D/Dan Brown - The Da Vinci Code.epub.txt
@@
4 Albert Camus (Justin O’Brien, translator) The Myth of Sisyphus [sisyphus] 1955 © the-eye.eu/public/Books/Bibliotik/2/2013(orig1942) Albert Camus - The Myth of Sisyphus[Transl Justin O’Brien]_Ral.epub.txt
@@
5 Lewis Carroll Alice’s Adventures in Wonderland [alice] 1865 PD the-eye.eu/public/Books/Bibliotik/A/Alice’s Adventures in Wonderlan - Lewis Carroll.epub.txt
@@
6 Ta-Nehisi Coates The Beautiful Struggle [struggle] 2009 © the-eye.eu/public/Books/Bibliotik/T/The Beautiful Struggle - Ta-Nehisi Coates.epub.txt
@@
7 Ta-Nehisi Coates We Were Eight Years in Power: An American Tragedy [wewere] 2017 © the-eye.eu/public/Books/Bibliotik/W/We Were Eight Years in Power - Ta-Nehisi Coates.epub.txt
@@
8 Ta-Nehisi Coates The Water Dancer: A Novel [waterdancer] 2019 © the-eye.eu/public/Books/Bibliotik/T/The Water Dancer - Ta-Nehisi Coates;.epub.txt
@@
9 Jon Cogburn Dungeons and Dragons and Philosophy [dungeons] 2012 © the-eye.eu/public/Books/Bibliotik/2/2012 Jon Cogburn - Dungeons and Dragons and Philosophy - Raiding the Temple of Wisdom_Rsnl.epub.txt
10 Junot Díaz Drown [drown] 1996 © the-eye.eu/public/Books/Bibliotik/J/Junot Diaz - Drown.epub.txt
@@
11 Junot Díaz The Brief Wondrous Life of Oscar Wao [oscarwao] 2007 © the-eye.eu/public/Books/Bibliotik/T/The Brief Wondrous Life of Oscar Wao.epub.txt
@@
12 Junot Díaz This Is How You Lose Her [loseher] 2012 © the-eye.eu/public/Books/Bibliotik/T/This Is How You Lose Her-Diaz.epub.txt
@@
13 Cory Doctorow Down and Out in the Magic Kingdom [downandout] 2004 $C$-BY-NC-SA the-eye.eu/public/Books/Bibliotik/D/Down and Out in the Magic Kingd - Cory Doctorow.epub.txt
@@
14 Carol Ann Duffy The World’s Wife [worldswife] 2001 © the-eye.eu/public/Books/Bibliotik/T/The World_s Wife - Carol Ann Duffy.epub.txt
@@
15 Jennifer Egan A Visit from the Goon Squad [goonsquad] 2010 © the-eye.eu/public/Books/Bibliotik/A/A\ Visit\ from\ the\ Goon\ Squad\ -\ Jennifer\ Egan.epub.txt
@@
16 Christopher Farnsworth The President’s Vampire [vampire] 2011 © the-eye.eu/public/Books/Bibliotik/T/The President’s Vampire - Christopher Farnsworth.epub.txt
@@
17 F. Scott Fitzgerald The Great Gatsby [gatsby] 1925 PD the-eye.eu/public/Books/Bibliotik/T/The Great Gatsby - F. Scott Fitzgerald.epub.txt
@@
18 Malcom Gladwell Blink: The Power of Thinking Without Thinking [blink] 2005 © the-eye.eu/public/Books/Bibliotik/B/Blink - Malcolm Gladwell.epub.txt
@@
19 Christopher Golden Dead Ringer [deadringer] 2016 © the-eye.eu/public/Books/Bibliotik/C/Christopher Golden - Dead Ringers [retail].epub.txt
20 Christopher Golden Ararat [ararat] 2017 © the-eye.eu/public/Books/Bibliotik/A/Ararat - Christopher Golden.epub.txt
@@
21 Andrew Sean Greer The Confessions of Max Tivoli [maxtivoli] 2005 © the-eye.eu/public/Books/Bibliotik/T/The Confessions of Max Tivoli - Andrew Sean Greer.epub.txt
@@
22 John Grisham Theodore Boone: The Fugitive [fugitive] 2015 © the-eye.eu/public/Books/Bibliotik/T/The Fugitive - John Grisham.epub.txt
@@
23 Mark Haddon The Curious Incident of the Dog in the Night-Time [curious] 2003 © the-eye.eu/public/Books/Bibliotik/T/The Curious Incident of the Dog in the Night-Time - Mark Haddon.epub(1).txt
24 Joseph Heller Catch-22 [catch22] 2018 © the-eye.eu/public/Books/Bibliotik/C/Catch-22 - Joseph Heller.epub.txt
@@
25 David Henry Hwang M. Butterfly [mbutterfly] 1988 © the-eye.eu/public/Books/Bibliotik/M/M. Butterfly - David Henry Hwang.epub.txt
26 Betty E.M. Jacobs All the Onions [onions] 1999 © the-eye.eu/public/Books/Bibliotik/A/All the Onions - Betty E. M. Jacobs.epub.txt
@@
27 James Joyce Ulysses [ulysses] 1922 PD the-eye.eu/public/Books/Bibliotik/U/Ulysses - James Joyce - Penguin Group - 2000.epub.txt
@@
28 Richard Kadrey Sandman Slim [sandmanslim] 2009 © the-eye.eu/public/Books/Bibliotik/S/Sandman Slim - Richard Kadrey.epub.txt
29 Matthew Klam Who Is Rich? [rich] 2017 © the-eye.eu/public/Books/Bibliotik/W/Who Is Rich_ - Matthew Klam.epub.txt
@@
30 Laura Lippman After I’m Gone [afterimgone] 2014 © the-eye.eu/public/Books/Bibliotik/A/After I’m Gone - Laura Lippman.epub.txt
31 Laura Lippman Sunburn [sunburn] 2018 © the-eye.eu/public/Books/Bibliotik/S/Sunburn - Laura Lippman.epub.txt
@@
32 George R.R. Martin A Game of Thrones [thrones] 1996 © the-eye.eu/public/Books/Bibliotik/A/A Game of Thrones - George R. R. Martin.epub.txt
33 Stephenie Meyer Twilight [twilight] 2005 © the-eye.eu/public/Books/Bibliotik/T/Twilight - Stephenie Meyer.epub.txt
34 Toni Morrison Beloved [beloved] 1987 © the-eye.eu/public/Books/Bibliotik/B/Beloved - Toni Morrison.epub.txt
@@
35 Yōko Ogawa The Memory Police [memory] 2019 © the-eye.eu/public/Books/Bibliotik/T/The Memory Police - Yoko Ogawa.epub.txt
@@
37 Philip Pullman The Subtle Knife [subtleknife] 1997 © the-eye.eu/public/Books/Bibliotik/P/Pullman - The Subtle Knife.epub.txt
38 Ramzi Rouighi The Making of a Mediterranean Emirate [emirate] 2011 © the-eye.eu/public/Books/Bibliotik/T/The Making of a Mediterranean E - Rouighi, Ramzi;.epub.txt
39 J.K. Rowling Harry Potter and the Sorcerer’s Stone [hp1] 1998 © the-eye.eu/public/Books/Bibliotik/H/Harry_Potter_ and_the_Sorcerers_Stone-Rowling.epub.txt
40 J.K. Rowling Harry Potter and the Goblet of Fire [hp4] 2000 © the-eye.eu/public/Books/Bibliotik/H/Harry_Potter_ and_the_Goblet_of_Fire-Rowling.epub.txt
41 J.D. Salinger The Catcher in the Rye [catcher] 1951 © the-eye.eu/public/Books/Bibliotik/T/The Catcher in the Rye - J. D. Salinger.epub.txt
42 Sheryl Sandberg Lean In [leanin] 2013 © the-eye.eu/public/Books/Bibliotik/L/Lean In - Sheryl Sandberg.epub.txt
43 Sarah Silverman The Bedwetter [bedwetter] 2010 © the-eye.eu/public/Books/Bibliotik/T/The Bedwetter - Sarah Silverman.epub.txt
44 Rachel Louise Snyder No Visible Bruises [novisible] 2019 © the-eye.eu/public/Books/Bibliotik/N/No Visible Bruises - Rachel Louise Snyder.epub.txt
45 Lysa TerKeurst Unglued [unglued] 2012 © the-eye.eu/public/Books/Bibliotik/L/Lysa TerKeurst, Unglued.epub.txt
46 Lysa TerKeurst Embraced [embraced] 2018 © the-eye.eu/public/Books/Bibliotik/E/Embraced - Lysa Terkeurst.epub.txt
47 J.R.R. Tolkien The Hobbit [hobbit] 1937 © the-eye.eu/public/Books/Bibliotik/T/The Hobbit (Houghton Mifflin Harcourt) (75th Anniversary Edition) [Epub] - J.R.R. Tolkien.epub.txt
48 Jacqueline Woodson Brown Girl Dreaming [browngirldreaming] 2014 © the-eye.eu/public/Books/Bibliotik/B/Brown Girl Dreaming - Jacqueline Woodson.epub.txt
49 Jacqueline Woodson Another Brooklyn [anotherbrooklyn] 2016 © the-eye.eu/public/Books/Bibliotik/A/Another Brooklyn - Jacqueline Woodson.epub.txt
50 Jonathan Zittrain The Future of the Internet and How to Stop It [zittrain] 2008 © the-eye.eu/public/Books/Bibliotik/J/Jonathan Zittrain - The Future of the Internet.epub.txt